OpenStudy (anonymous):

sin(theta) + cos(theta) = square root of 2

OpenStudy (anonymous):

i would guess \[\frac{\pi}{4}\]

OpenStudy (anonymous):

since both sine and cosine give \[\sin(\frac{\pi}{4})=\cos(\frac{\pi}{4})=\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

how do we know theta = pi/4?

OpenStudy (anonymous):

because i know that both sine and cosine are \[\frac{\sqrt{2}}{2}\] there so when you add them you will get \[\frac{\sqrt{2}}{2}+\frac{\sqrt{2}}{2}=\sqrt{2}\]

OpenStudy (anonymous):

Okay make sense. Thank you so much!

OpenStudy (anonymous):

yw